Suggest me in math how to calculate(adding,multification,etc) fast, so please tell me some expensive tricks ,
THANK U

Grade:12

1 Answers

Vasantha Kumari
9 years ago

Addition:

1862+1928

1862 = 1000 + 800 + 60 + 2

1928 = 1000 + 900 +20 + 8

= 2000 + 1700 + 80 +10 = 3790

Subtraction:

For subtracting 36 from 62 split as shown

6 / 2

–3 / 6

-----------

3 / (–4)

Here the answer is negative we are in need to add 10 with the second digit and subtract 1 with the first digit obtained after subraction.

3 /(–4)

(3–1)/(10–4)

----------

2/6------Answer is 26

Multiplication:

Use the nearest base for quick multiplication.

For example to multiply 8 ´ 7

So, 8 is two below 10 and 7 is three below 10, so

8 2

´ |

7 3

--------

5 6 answer

-------

For squaring numbers:

For squaring 75, the last part is always 25 and 7 has to be multiplied to a number more than it…i.e., 8 which is equal to 56.

7 52 = 56 25

----- 7 ´ 8 = 56

Thus answer is 5625

Hope I have given expensive tricks as asked by you,


Thanks & Regards,

Vasantha Sivaraj,

askIITians faculty